OpenStudy (anonymous):

If the point (1, 5) lies on a circle whose center is at the origin, what is the equation of the circle? A.) x(squared) + y(squared) = 26 B.) x(squared) + y(squared) = 16 C.) x(squared) + y(squared) = 13 D.) x(squared) + y(squared) = 11

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

the anwer is C I THINK

OpenStudy (anonymous):

The radius is the distance from the center to the circle, so use the distance formula again to find the radius. r = distance between (0,0) and (1,5)

OpenStudy (dumbcow):

the equation for a circle actually comes directly from the distance formula

OpenStudy (anonymous):

I got b, is that right?

OpenStudy (anonymous):

NO ITS c

OpenStudy (anonymous):

no.

OpenStudy (dumbcow):

hmmm... \[x^2 + y^2 = r^2\] (x,y) = (1,5) \[1^2 + 5^2 = 1+25=26\]

OpenStudy (anonymous):

ITS CCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCC